# Testing stimulated emission photon directions

### v8

@@ Abstract
While naively laser only causes excitation of external target, e.g. Rabi cycle, STED microscopy or backward ASE in optical amplifiers demonstrate it can also cause its deexitation, however, under uncommon condition of being prepared as excited. These two causalities are governed by absorption-stimulated emission pair of equations, and swap places in perspective of CPT symmetry, however, it means photon direction of stimulated emission should be opposite to usually assumed. This article discusses various arguments and proposes simple direct tests to experimentally verify existence of such backward photon trajectories, complementing consequent forward textbook trajectories. There are already indirect experimental arguments, like backward ASE, lack of Rabi cycle for free electron laser, or recent ”negative time evidence” experiment in the news. Depending on the result: it should lead to new applications like 2WQC more symmetric quantum computers, or macroscopic violation of CPT symmetry, so far probably tested only in microscopic settings. Also proposed forward/backward/bidirectional photon sources are discussed.
@@ I Introduction
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [4] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature. It has dozens of successful experimental confirmations [5], however, it seems all tests so far were microscopic - come from observations of single particles. Conscious macroscopic test was not found in literature, important especially due to T time symmetry which might lead to new applications - some are proposed and discussed here. For example through system response before impulse, like recently experimentally observed by Steinberg group in the news presented in Fig. 1 [1]111 www.scientificamerican.com/article/evidence-of-negative-time-found-….
@@
Multiple updates of this article were mainly motivated by finding additional indirect experimental evidence in the literature (e.g. for Fig. 1, 3, 5), transparency with textbook view as consequent, convenient optical amplifier realization, and adding multiple potential applications. Also, QInter of our 2WQC group (https://www.qaif.org/2wqc): Vysakh Saji has performed the proposed experiment, obtaining looking confirming results, but without followup. Recently we have started local tests (with Tomasz Urbańczyk and Tomasz Kawalec), which lead to observation/difficulty that nearly all available laboratory lasers have built-in optical isolator, what seems one of reasons backward ASE properties are practically unknown, also confirmed by discussion with multiple photonics and fiber optics experts. This article should be updated with further tests.
@@ II-B Alternative STED-like tests
Center: STED microscopy (image source) showing that, at least for diode lasers, both equations act on external target, hence we can call it bidirectional beam. We usually focus on absorption by external target, because stimulated emission requires its initial excitation, while without such preparation usually $N_{2}/N\approx 0$. In Rabi cycle laser periodically causes excitation and deexitation of target, suggesting requirement of bidirectional beam. Its experimental support is lack of Rabi oscillations for free electron laser [17] - leading to $N_{2}\approx N$ nearly complete excitation in Autler-Townes effect (dynamic Stark), instead of $N_{2}\approx N/2$ for damped Rabi cycle.
@@ III-C Rabi cycle as coupling, its asymmetrization
Such bidirectional energy exchange for e.g. laser-atoms coupling is mainly realized by optical photons, hence we can view standard e.g. diode laser as bidirectional beam - causing both excitation of target, but also deexcitation - e.g. in Rabi cycle, STED microscopy, or backward ASE of EDFA.
Being able to introduce asymmetry in this coupling could lead to new applications. For example free electron laser (FEL) does not have such resonator coupling, hence instead of damped Rabi [27] leading to $N_{2}\to N/2$, FEL experiments [17] show $N_{2}\to N$ only excitation with unidirectional: forward beam - photon exchange only in one direction, acting only with absorption equation on the target. CPT symmetry suggests that looking at such synchrotron radiation beam from CPT perspective, photon beam created by reversed electron trajectory should act as backward beam - causing only target deexctiation, acting only with stimulated emission equation.
@@
Center: vacuum Rabi cycle [21] - observed oscillations between optical cavity resonator, and atom as tiny resonator - of energy between being localized in atom (high energy density), and delocalized standing EM wave (low energy density) - having different EM fields, like localized laser vs delocalized radio communication. Laser can enforce Rabi oscillations, periodically causing target excitation and deexcitation - allowing to view these causalities as results of coupling of two resonators, in contrast e.g. to free electron laser without resonator - hence only observing target excitations there [17]: $(N_{2}\to N)$ instead of damped Rabi $(N_{2}\to N/2)$. Being able to add some asymmetry to such energy transfer caused by coupling, as discussed further can lead to many applications - e.g. optical isolator brings hope for that: designed to block localized optical photons in one direction, what not necessarily has to be true for delocalized standing wave for coupling - having lower energy density.
Bottom: turns out nearly all lasers have built-in optical isolator, what might prevent required Rabi bidirectional energy exchange like for FEL. Searching literature, while for microwaves Rabi cycle has straightforward realization, for lasers there are used tricks like added resonator [29], or Mach-Zehnder type setting ([30, 31]), requiring analysis like in ”Asking photons where they have been” [32]. Just removing this isolator might allow for simpler realizations.
Center: therefore, Rabi cycle also needs hydrodynamical analog, tank of liquid with standing wave seems analogous to laser resonator. Being semi-transparent/semi-open in one direction (also in Fig. 7), it should cause left-right swimming of a probe particle outside, through exchange of momentum acting in both directions, using positive/negative pressure. In CPT perspective it looks the same, also with energy/momentum exchange in both directions.
Bottom: while the above bidirectional source is resonator extending to external targets, coupling with them through standing wave e.g. for Rabi oscillations, there are also unidirectional sources. In this analogy, synchrotron radiation is kind of water stream allowing only unidirectional momentum transfer. Optical isolator analog could be some unidirectional valve, analog of ring laser: a pump enforcing unidirectional circulation - with reversed circulation in CPT perspective. Radiation pressure is a vector: can be also negative [26], suggesting to build more symmetric and powerful two-way microfluidic, photonic computers as in Fig. 11.
@@ IV Some potential new applications
  • Forward beam, e.g. from synchrotron radiation, could allow for state preparation of quantum computers to nearly completely excited $N_{2}\approx N$ as in [17], backward beam for its CPT analog (postparation) for 2WQC [10] as in Fig. 11, in theory allowing for up to exponential speedup and better error correction ([34, 35]).
@@
  • Backward beam in theory should be transparent for $N_{2}=0$ medium, what could allow for much better penetration than forward (transparent for $N_{1}=0$) - e.g. for radiotherapy as in Fig. 12, backward camera as in Fig. 10 and various scanners e.g. CAT for 3D map of $N_{2}$ atoms/molecules as in Fig. 13, communication through obstacles, etc.
@@
  • Backward beam to stimulate specific chemical processes: releasing emission spectrum photons of chosen energy (and polarization), e.g. for chemical/technological processes, energy production, or medical therapies as in Fig. 12 - e.g. locally reduce lifetime of NADH molecule to starve cancer tissue and release this energy as thermal, or inhibit biochemical processes by amplifying deexcitation of crucial molecules or injected contrast medium, for example with multiple rotating beams intersecting in cancer tissue, especially if better transparency is confirmed.
@@
  • Backward camera as in Fig. 10 maybe exploiting transparency and/or opposite delay e.g. for astronomical observations (for example pulsars could produce forward+backward synchrotron radiation) - focused on target’s backward beam (or transmission of added backward beam), while standard camera is focused on its forward beam. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited approximately 3-level medium like dye or erbium, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency). Pairing with backward beam in line, could allow for various scanning (e.g. medical, material science, geological) allowing (emission) CAT scanning for 3D maps of $N_{2}$ excited atoms/molecules between them as in Fig. 10, 13.
    Figure 9: While most laboratory lasers have built-in (forward) optical isolator - can be referred as forward laser (only pushing photons), especially for Rabi cycle as in Fig. 7 it is worth to distinguish (bidirectional or) just laser having removed optical isolator. For various mentioned applications it is also worth to consider backward laser as CPT transformed forward laser: with reversed optical isolator, should only pull photons. Polarization independent optical isolators usually change angle for the opposite photon direction, allowing to add mirrors inside to reflect back such photons (e.g. extending laser resonator toward one time direction) - mirrored backward laser with isolator containing such mirrors should be much more energy efficient, might also improve forward laser as having lower energy loss toward $-t$. Angle, polarization, path for to reflect such reverse photons would need optimization, similar effect could be realized with circulator, ideally would be getting such effect with metamaterials.
    Figure 10: Top: Backward camera/detector observing backward ASE caused by the target (or added backward laser), especially interesting if confirmed better transparency and/or opposite delay sign. Detector can be e.g. based on (multi-mode) fiber with circulators as in [7] and Fig. 3. For camera there can be used semitransparent mirror and CCD/CMOS - both need some continuously pumped medium (maybe from some display technology). Alternatively one could monitor spontaneous emission of the pumped medium, e.g. with camera behind. Pairing backward detector/camera with pointing backward laser could allow for transmission scanning for up to 3D maps of excited molecules between (e.g. medical, material science, geological scanning) - like in standard transmission scanning, but of emission spectrum instead of absorption.
@@
  • Backward-lighted camera as in Fig. 10: with added backward beam for observations actively increasing deexcitation rate of excited objects, e.g. for thermal imaging.
    Figure 11: Original motivation: (example of) two-way quantum computer (2WQC) [10] (and its classical hydrodynamical analog) as adding a process which in CPT perspective becomes the original state preparation process. For photonic QC state preparation is usually performed by laser impulse, suggesting to just add ’reversed’ impulse: delayed by propagation time, and behind backward placed optical isolator - this way from CPT perspective photon trajectories should be reversed, like improving two-way flow control in hydrodynamical analog. In comparison to standard 1WQC, in theory 2WQC allows for up to exponential speedup e.g. to solve NP problems, also better stability and error correction [34, 35]. Such classical 2WCC photonic computer on continuous beam or impulses seems also worth considering.
    Figure 12: If being able to separate absorption and stimulated emission equations, for example with CPT analog of synchrotron radiation: beam resulting from reversed electron trajectories as in Fig. 6, or maybe standard laser with backward placed optical isolator as in Fig. 9, one of many new applications could be radiotherapy trying to locally amplify emission of photons of chosen energy and polarization. For example to try to locally increase degradation of energy carrying molecules like NADH (nicotinamide adenine dinucleotide) crucial for cell energy metabolism, or other autofluorescent molecules with emission spectrum (lists e.g. in [37, 38, 39]) - to starve cancer tissue, and heat up by releasing this energy. Intersecting multiple such beams, or e.g. one beam spread in 1D by two cylindrical lenses, additionally rotating them for fixed focus intersection point, we could make the (usually nonlinear) effect there orders of magnitude stronger then for intermediate tissues, to make it negligible outside of targeted tissue. Different issues could be also approached this way, e.g. contrast agent as in photodynamic therapy [40], or trying to disable toxic molecules (like $\sim$1280nm for singlet oxygen [41]), inhibit chosen biochemical pathways by stimulating deexctiation of crucial molecules e.g. of pathogens, etc.
  • Backward beam for 3D printing/reversed lithography: excite material and precisely stimulate deexcitation - controlling position, time, polarization, also frequency to choose one of multiple possible transitions e.g. of high energy to locally reduce temperature (cooling), of energy regained thermally.
  • Synchrotron sources usually have much higher photon energy, its backward beam could e.g. increase probability of selected nuclear decays/transitions: producing photons of chosen energy, polarization. Maybe also for other particle events: which CPT transform requires such photons incoming from e.g. synchrotron.
Figure 13: Simplified Fig. 10 diagram for measurement of (integrated over light path) absorption coefficient, which can be extended e.g. to CAT scanner creating 3D maps of absorption coefficient from 2D projections in multiple angles, reconstructed through Radon transform [42]. Above CPT symmetric setting suggests how to analogously measure emission coefficients, what could be extended to its CAT scanning. For medical applications e.g. of autoluminescent molecules (gaining energy from cell chemistry) to make 3D maps of e.g. tryptophan ($\sim$340nm), collagen ($\sim$360nm), elastin ($\sim$410nm), NADH ($\sim$460nm), flavins ($\sim$525nm) [37], or some contrast agent - especially if possible to separate, and indeed being blocked only by molecules emitting in such chosen frequency (better transparency). Using optimized multiple frequencies, could allow for noninvasive detailed (high spatial and temporal resolution) 3D maps of tissues with detected pathologies like cancer. The source could be laser with reversed optical isolator, or ring laser, or synchrotron radiation. The detector could be backward camera as in Fig. 10: pumped surface (e.g. screen) with monitored population, or adding e.g. half-silvered mirror and standard detector.

### v9

@@ Abstract
While naively laser only causes excitation of external target, e.g. Rabi cycle, STED microscopy or backward ASE in optical amplifiers demonstrate it can also cause its deexitation, however, under uncommon condition of being prepared as excited. These two causalities are governed by absorption-stimulated emission pair of equations, and swap places in perspective of CPT symmetry, however, it means photon direction of stimulated emission should be opposite to usually assumed. This article discusses various arguments and proposes simple direct tests to experimentally verify existence of such backward photon trajectories, complementing consequent forward textbook trajectories. Depending on the results, it could lead to many proposed applications like 2WQC more symmetric quantum computers, or if unsuccessful to macroscopic violation of CPT symmetry, so far probably tested only in microscopic settings.
@@ I Introduction
CPT symmetry (of charge conjugation (C) + parity transformation (P) + time reversal (T)) is shown [4] to be necessary for local Lorentz invariant quantum field theories, hence is generally believed to be satisfied by our nature. It has dozens of successful experimental confirmations [5], however, it seems all tests so far were microscopic - come from observations of single particles. Conscious macroscopic test was not found in literature, important especially due to T time symmetry which might lead to new applications - some are proposed and discussed here. For example through system response before impulse, like recently experimentally observed by Steinberg group as in Fig. 1 [1]111 www.scientificamerican.com/article/evidence-of-negative-time-found-….
@@
Multiple updates of this article were mainly motivated by finding additional indirect experimental evidence in the literature (e.g. for Fig. 1, 3, 5), transparency with textbook view as consequent, convenient optical amplifier realization, and adding multiple potential applications. Also, QInter of our 2WQC group (https://www.qaif.org/2wqc): Vysakh Saji has performed the proposed experiment, obtaining looking confirming results, but without followup. Recently we have started local tests (with Tomasz Urbańczyk and Tomasz Kawalec), which lead to observation/difficulty that most of available laboratory lasers have built-in optical isolator, what seems one of reasons backward ASE properties are practically unknown, also confirmed by discussion with multiple photonics and fiber optics experts. Further tests seem to show pulling of photons as required by CPT symmetry and discussed here - planned to be developed and described in a separate article, also to update this article.
@@ II-B Alternative STED-like tests
Center: STED microscopy (image source) showing that, at least for diode lasers, both equations act on external target, hence we can call it bidirectional beam. We usually focus on absorption by external target, because stimulated emission requires its initial excitation, while without such preparation usually $N_{2}/N\approx 0$.
@@
Center: vacuum Rabi cycle [23] - observed oscillations between optical cavity resonator, and atom as tiny resonator - of energy between being localized in atom (high energy density), and delocalized standing EM wave (low energy density) - having different EM fields, like localized laser vs delocalized radio communication. Laser can enforce Rabi oscillations, periodically causing target excitation and deexcitation - allowing to view these causalities as results of coupling of two resonators. Being able to add some asymmetry to such energy transfer caused by coupling, as discussed further can lead to many applications - e.g. optical isolator brings hope for that: designed to block localized optical photons in one direction, what not necessarily has to be true for delocalized standing wave for coupling - having lower energy density.
Bottom: turns out nearly all lasers have built-in optical isolator, what might prevent required Rabi bidirectional energy exchange like for FEL. Searching literature, while for microwaves Rabi cycle has straightforward realization, for lasers there are used tricks like added resonator [24], or Mach-Zehnder type setting ([25, 26]), requiring analysis like in ”Asking photons where they have been” [27]. Just removing this isolator might allow for Rabi simpler realizations. The Mach-Zehnder setting also suggests (to be tested) how to get backward beam having laser with built-in optical isolator: by just adding splitter/coupler.
Center: therefore, also Rabi cycle also needs hydrodynamical analog, tank of liquid with standing wave seems analogous to laser resonator. Being semi-transparent/semi-open in one direction (also in Fig. 7), it should cause left-right swimming of a probe particle outside, through exchange of momentum acting in both directions, using positive/negative pressure. In CPT perspective it looks the same, also with energy/momentum exchange in both directions.
Bottom: while the above bidirectional source is resonator extending to external targets, coupling with them through standing wave e.g. for Rabi oscillations, there are also unidirectional sources. In this analogy, synchrotron radiation is kind of water stream allowing only unidirectional momentum transfer. Optical isolator analog could be some unidirectional valve, analog of ring laser: a pump enforcing unidirectional circulation - with reversed circulation in CPT perspective. Radiation pressure is a vector: can be also negative [29, 30], suggesting to build more symmetric and powerful two-way microfluidic, photonic computers as in Fig. 10.
@@ III-C Rabi cycle as coupling, its asymmetrization
Figure 9: While most laboratory lasers have built-in (forward) optical isolator - can be referred as forward laser (only pushing photons), especially for Rabi cycle as in Fig. 7 it is worth to distinguish (bidirectional or) just laser having removed optical isolator. For various mentioned applications it is also worth to consider backward laser as CPT transformed forward laser: with reversed optical isolator, should only pull photons with negative radiation pressure. Polarization independent optical isolators usually change angle for the opposite photon direction, allowing to add mirrors inside to reflect back such photons (e.g. extending laser resonator toward one time direction) - mirrored backward laser with isolator containing such mirrors should be much more energy efficient, might also improve forward laser as having lower energy loss toward $-t$. Angle, polarization, path for to reflect such reverse photons would need optimization, similar effect could be realized with circulator, ideally would be getting such effect with metamaterials.
Figure 10: Original motivation: (example of) two-way quantum computer (2WQC) [10] (and its classical hydrodynamical analog) as adding a process which in CPT perspective becomes the original state preparation process. For photonic QC state preparation is usually performed by laser impulse, suggesting to just add ’reversed’ impulse: delayed by propagation time, and behind backward placed optical isolator - this way from CPT perspective photon trajectories should be reversed, like improving two-way flow control in hydrodynamical analog. In comparison to standard 1WQC, in theory 2WQC allows for up to exponential speedup e.g. to solve postBQP [33], NP problems, also better stability and error correction [34, 35]. Such classical 2WCC photonic computer on continuous beam or impulses seems also worth considering.
Figure 11: Top: Backward camera/detector observing backward ASE caused by the target (or added backward laser), especially interesting if confirmed better transparency and/or opposite delay sign. Detector can be e.g. based on (multi-mode) fiber with couplers/circulators as in [7] and Fig. 3. For camera there can be used semitransparent mirror and CCD/CMOS - both need some continuously pumped medium (maybe from some display technology). Alternatively one could monitor spontaneous emission of the pumped medium, e.g. with camera behind. There could be added optical isolator to remove incoming photons. As in Fig. 13, pairing backward detector/camera with pointing backward laser could allow for transmission scanning for up to 3D maps of excited molecules between (e.g. medical, material science, geological scanning) - like in standard transmission scanning, but of emission spectrum instead of absorption.
Figure 12: If being able to separate absorption and stimulated emission equations, for example with CPT analog of synchrotron radiation: beam resulting from reversed electron trajectories as in Fig. 6, or maybe standard laser with backward placed optical isolator as in Fig. 9, one of many new applications could be radiotherapy trying to locally amplify emission of photons of chosen energy and polarization. For example to try to locally increase degradation of energy carrying molecules like NADH (nicotinamide adenine dinucleotide) crucial for cell energy metabolism - to starve cancer tissue, and heat up by releasing this energy. Or other autofluorescent molecules with emission spectrum (lists e.g. in [36, 37, 38]) to locally influence metabolism. Intersecting multiple such beams, or e.g. one beam spread in 1D by two cylindrical lenses, additionally rotating them for fixed focus intersection point, we could make the (usually nonlinear) effect there orders of magnitude stronger then for intermediate tissues, to make it negligible outside of targeted tissue. Different issues could be also approached this way, e.g. applying contrast agent as in photodynamic therapy [39], or trying to disable toxic molecules (like $\sim$1280nm for singlet oxygen [40]), inhibit chosen biochemical pathways by stimulating deexctiation of crucial molecules e.g. of pathogens, etc.
Such bidirectional energy exchange for e.g. laser-atoms coupling is mainly realized by optical photons, hence we can view standard e.g. diode laser as bidirectional beam - causing both excitation of target, but also deexcitation - e.g. in Rabi cycle, STED microscopy, or backward ASE of EDFA.s
Being able to introduce asymmetry in this coupling could lead to new applications. For example free electron laser (FEL) does not have such resonator coupling, CPT symmetry suggests that looking at such synchrotron radiation beam from CPT perspective, photon beam created by reversed electron trajectory should act as backward beam - causing only target deexctiation, acting only with stimulated emission equation.
@@ IV Some potential new applications
Figure 13: Simplified diagram for measurement of (integrated over light path) absorption coefficient, which can be extended to 3D mapping in CT scanner (computer tomography) from 2D projections in multiple angles, reconstructed through Radon transform [41]. Above CPT symmetric setting suggests how to analogously measure emission coefficients, what could be extended to its 2D projections, videos, and finally CT scanning for 3D maps. For medical applications e.g. of autoluminescent molecules (gaining energy from cell chemistry) to make 3D maps of e.g. tryptophan ($\sim$340nm), collagen ($\sim$360nm), elastin ($\sim$410nm), NADH ($\sim$460nm), flavins ($\sim$525nm) [36], or some contrast agent - especially if possible to separate, and indeed being blocked only by molecules emitting in the used frequency (better transparency). Using optimized multiple frequencies, could allow for noninvasive detailed (high spatial and temporal resolution) 2D color photos/videos, or 3D maps of tissues with detected pathologies like cancer - multispectral, without ionizing radiation. The source could be laser with reversed optical isolator, or ring laser, or synchrotron radiation. The detector could be backward camera as in Fig. 11: pumped surface (e.g. screen) with monitored population, or adding e.g. half-silvered mirror and standard detector.
  • Forward beam, e.g. from synchrotron radiation, could allow for state preparation of quantum computers to nearly completely excited $N_{2}\approx N$ as in [42], backward beam for its CPT analog (postparation) for 2WQC [10] as in Fig. 10, in theory allowing for up to exponential speedup (solve postBQP [33]) and better error correction ([34, 35]).
@@
  • Backward beam in theory should be transparent for $N_{2}=0$ medium, what could allow for much better penetration than forward (transparent for $N_{1}=0$) - e.g. for radiotherapy as in Fig. 12, backward camera as in Fig. 11 and various scanners e.g. CT for 3D map of $N_{2}$ atoms/molecules as in Fig. 13, communication through obstacles, etc.
@@
  • Backward beam to stimulate specific atomic, chemical or nuclear processes: releasing emission spectrum photons of chosen energy (and polarization), e.g. for chemical/technological processes, energy production, or medical therapies as in Fig. 12 - e.g. locally reduce lifetime of NADH molecule to starve cancer tissue and release this energy as thermal, or inhibit biochemical processes by amplifying deexcitation of crucial molecules or injected contrast medium, for example with multiple rotating beams intersecting in cancer tissue, especially if better transparency is confirmed.
@@
  • Backward camera as in Fig. 11 maybe exploiting transparency and/or opposite delay e.g. for astronomical observations (of e.g. synchrotron radiation of pulsars or our Sun [43] as in Fig. 14) - focused on target’s backward beam (or transmission of added backward beam), while standard (forward) camera is focused on its forward beam. It would replace CMOS/CCD matrix behind the lens, with matrix/surface of continuously excited e.g. approximately 3-level medium like dye or erbium, plus some monitoring of its population map - e.g. through spontaneous emission, or EIT (electromagnetically induced transparency), or half-silvered mirror and CMOS/CCD. Pairing with backward beam could allow for various scanning (e.g. medical, material science, geological) allowing (emission) CT scanning for 3D maps of $N_{2}$ excited atoms/molecules between them as in Fig. 13.
    Figure 14: Potential application of proposed backward camera to obtain complementary information of astronomical objects, especially interesting if delay sign and transparency are confirmed. Many of such objects e.g. pulsars (image source) have accelerating charged particles, what leads to synchrotron radiation - e.g. causing excitation in cooled detector of standard telescopes. In CPT perspective of this scenario, there is still accelerating charge (in opposite direction) - should produce photons, causing excitation now toward our negative time, which is deexcitation toward positive time. Therefore, such process required by CPT symmetry would need excited target - what is rarely true, what might be the reason for not observing such effect, could lead to worth searching consequences like deexcited halos around such pulsars. To have a chance to observe such effect, we would need to pump the sensor/detector (instead of standard cooling), again monitoring its population, or maybe using semi-transparent mirror as in Fig. 11, maybe also include some optical isolator to allow only outgoing photons. If successful, such telescope should see (reversed) synchrotron, but not thermal radiation - e.g. allowing to separate them when combining with standard picture, also for observations/diagnostics/early warnings for our Sun [43].
@@
  • Backward-lighted camera as in Fig. 11: with added backward beam for observations actively increasing deexcitation rate of excited objects, e.g. for thermal imaging, of contrast agent.
  • Backward beam for 3D printing/reversed photolithography: excite material and precisely stimulate deexcitation (can be in 3D as for radiotherapy) - controlling position, time, frequency, also polarization to choose one of multiple possible transitions e.g. of high energy for local cooling.
  • Backward beam for optical pulling, directly acting with negative radiation pressure.
  • Synchrotron sources usually have much higher photon energy, its backward beam could e.g. increase probability of selected nuclear decays/transitions: producing photons of chosen energy. Maybe also increasing probability of other particle events/Feynman diagrams: which CPT transform requires such photons incoming from e.g. synchrotron.